OpenStudy (anonymous):

Question. Consider the following. f(x) = x − sqrtx((x − 1)) Estimate the limit of f(x)as x approaches infinity and as x approaches negative infinity. Need help please:)

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}(x-\sqrt{x(x-1)}) =\lim_{x \rightarrow \infty}(x-\sqrt{x(x-1)}) \cdot \frac{x+\sqrt{x(x-1)}}{x+\sqrt{x(x-1)}} \\ =\lim_{x \rightarrow \infty}\frac{x^2-x(x-1)}{x+\sqrt{x(x-1)}}=\lim_{x \rightarrow \infty}\frac{x^2-x^2+x}{x+\sqrt{x(x-1)}}=\lim_{x \rightarrow \infty}\frac{x}{x+\sqrt{x^2-x}}\] Try dividing top and bottom by sqrt(x^2) sqrt(x^2)=x since x->infty (so x is positive) -- for the problem you would sqrt(x^2)=-x since x->-infty (so x is negative)

OpenStudy (freckles):

However since it says estimate I wonder if would rather you use a numerical approach instead of an algebraic one.

OpenStudy (anonymous):

I think numerical

OpenStudy (anonymous):

x-> negative infitinity

OpenStudy (freckles):

Well see if the x-sqrt(x(x-1)) 's outputs are getting to some number as you enter in like x=-20, x=-100, x=-1000

OpenStudy (freckles):

\[-20-\sqrt{-20(-20-1)}=? \\ -100-\sqrt{-100(-100-1)}=? \\ -1000-\sqrt{-1000(-1000-1)}=?\]

OpenStudy (anonymous):

These are my outputs -10^6 = -2000000.5000 -10^4 = -20000.5000 -10^2=-200.4988 10^0 = 1 10^2 = .5013 10^4 = .5000 10^6 = .5000

OpenStudy (anonymous):

I don't know if these would help.

OpenStudy (freckles):

Ok notice for x->-infty, we have the y's aren't approaching some number But for x->infty, we have the y's are approaching some number (that number being 1/2 or.5)

OpenStudy (freckles):

you could say as x->-infty, y->-infty

OpenStudy (freckles):

Because the y-values are getting negative huge

OpenStudy (anonymous):

ahh I see now:)

OpenStudy (freckles):

Also I would write your one post up there as These are my outputs Let f(x)=x-sqrt(x(x-1)) f(-10^6) = -2000000.5000 f(-10^4) = -20000.5000 f(-10^2)=-200.4988 f(10^2) = .5013 f(10^4) = .5000 f(10^6) = .5000